OpenStudy (anonymous):

A constant voltage is applied across a circuit. If the resistance in the circuit is doubled, what is the effect on the power dissipated by the circuit? The power dissipated is reduced by a factor of 4. The power dissipated is quadrupled. The power dissipated is doubled. The power dissipated is reduced by a factor of 2. The power dissipated remains constant.

OpenStudy (***[isuru]***):

@hartnn

hartnn (hartnn):

power dissipated = i*V = V * (V/R) = V^2/ R since V is constant, power dissipated is inversely proportional to R

hartnn (hartnn):

so when R doubles, Power dissipated halves

OpenStudy (anonymous):

ohh okay, so it would reduced by a factor of 2...thank you

hartnn (hartnn):

yes! welcome ^_^